what is 5x-18=4x+7
Answer by mananth(16949)

5x-18=4x+7
eliminate 4x by adding -4x
5x-4x-18=4x-4x+7
x-18=7
eliminate 18 by adding 18
x-18+18=7+18
x=25
Ananth